OpenStudy (anonymous):

8. Multiply: 9(6x +7) I need help ty

OpenStudy (kuoministers):

2(x+3) would be 2x + 6 because you would multiply 2 to x and 3

OpenStudy (anonymous):

\[9\times 6x+9\times 7\] is a start

OpenStudy (kuoministers):

can you work it out now or need more help?

OpenStudy (anonymous):

Need more help? I get the 9 x 6x and the 9x7 so it would be 54x+ 63 then I think it would be 117x?

OpenStudy (kuoministers):

You cannot add 54x to 63 because the bases number is NOT the same

OpenStudy (kuoministers):

in this case x is the base 4x + 83x = 87x

OpenStudy (kuoministers):

so in this case 54x+ 63 is the final answer do you understand it now?

OpenStudy (anonymous):

can u help me out by working it out step by step on the drawing board I would appreciate it.

OpenStudy (anonymous):

if x was $1000 then 54x would be $54000 but $54000 + 63 is not $117000

OpenStudy (kuoministers):